# Data-driven Implementations of Various Generalizations of Balanced Truncation

### v2

@@ 3.5 ADI/RADI-based Non-intrusive Implementation of BST
2: Set $D=G(\infty)$ and $R_{\text{S}}=DD^{T}$.
3: Set $S_{v}$, $L_{v}$, $S_{w}$, and $L_{w}$ as in (6).
4: Compute $Q_{v}$ and $P_{w}$ by solving the Lyapunov equations (9) and (10).
@@
6: Set $\hat{B}_{\text{S}}=T_{w}^{*}\big((\hat{W}^{*}\hat{V})Q_{v}^{-1}(C\hat{V})^{*}+(\hat{W}^{*}B)D^{T}\big)R_{\text{S}}^{-\frac{1}{2}}$.
7: Update $P_{w}$ by solving the Sylvester equation (77).
8: Decompose $Q_{v}^{-1}$ and $P_{w}^{-1}$ as $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$.
@@
10: Set $W_{r}=T_{w}\hat{L}_{q}\tilde{U}_{1}\tilde{\Sigma}_{r}^{-1/2}$ and $V_{r}=\hat{L}_{p}\tilde{V}_{1}\tilde{\Sigma}_{r}^{-1/2}$.
11: Compute $\hat{A}$, $\hat{B}$, and $\hat{C}$ from (28).
4 Practical Considerations in Implementation
The theoretical foundations of the ADI [16] and RADI [4] methods are well established. Their performance in approximating Lyapunov and Riccati equations is also well documented. However, the implementation presented in this paper involves inverses $Q_{v}^{-1}$ and $P_{w}^{-1}$, which can cause numerical problems as the number of samples used for implementation increases. This is because the solution of Lyapunov equations is generally low-rank, enabling low-rank approximation via the low-rank ADI method [16] in the first place. Thus, as the number of samples used for implementation increases, the matrices $Q_{v}$ and $P_{w}$ start losing rank. Note that these matrices do not appear in the original algorithms. In data-driven modeling, one typically uses as many samples as available to ensure maximum accuracy. As discussed in [2], the inverses $Q_{v}^{-1}$ and $P_{w}^{-1}$ in DD-ADI-BT can be computed conveniently when the interpolation points are lightly damped, i.e.,
$\sigma_{i}=\Bigg(\frac{\zeta_{i,\sigma}|\omega_{i,\sigma}|}{\sqrt{1-\zeta_{i,\sigma}^{2}}}\Bigg)+j\omega_{i,\sigma}\quad\textnormal{and}\quad\mu_{i}=\Bigg(\frac{\zeta_{i,\mu}|\omega_{i,\mu}|}{\sqrt{1-\zeta_{i,\mu}^{2}}}\Bigg)+j\omega_{i,\mu},$ (78)
wherein $\zeta_{i,\sigma}\ll 1$ and $\zeta_{i,\mu}\ll 1$. If we select the damping coefficients $\zeta_{i,\sigma}\approx 0$ and $\zeta_{i,\mu}\approx 0$, this means we are sampling $G(s)$ in close proximity to the $j\omega$-axis, which is also practically reasonable for capturing the frequency response of $G(s)$. When the interpolation points are lightly damped (i.e., $\zeta_{i,\sigma}\ll 1$ and $\zeta_{i,\mu}\ll 1$), the respective matrices $Q_{v}$, $P_{w}$, $Q_{v}^{-1}$, $P_{w}^{-1}$, $T_{v}$, and $T_{w}$ become diagonally dominant due to the modal form of the matrices $S_{v}$ and $S_{w}$ that appear in their associated Lyapunov, Sylvester, and Riccati equations. If $\zeta_{i,\sigma}\approx 0$ and $\zeta_{i,\mu}\approx 0$, the off-diagonal elements are nearly zero. For details and analytical formulas for computing the diagonal elements of the Lyapunov, Sylvester, and Riccati equations in this case, see [40]. Essentially, this means that each diagonal element of $Q_{v}^{-1}$ and $P_{w}^{-1}$ depends solely on its respective pair $(\sigma_{i},G(\sigma_{i}))$ or $(\mu_{i},G(\mu_{i}))$, while the dependence on other samples or sampling points becomes negligible.
Assuming that the interpolation points are lightly damped, the expressions for the diagonal entries of $Q_{v}^{-1}$ and $P_{w}^{-1}$ can be derived using the analytical formulas in [40] as follows:
For DD-LQG-BT:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\sigma_{1})[I_{m}+H^{*}(\sigma_{1})H(\sigma_{1})]^{-1},\cdots,$
$\hskip 113.81102pt2\text{Real}(\sigma_{v})[I_{m}+H^{*}(\sigma_{v})H(\sigma_{v})]^{-1}\big),$ (79)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\mu_{1})[I_{p}+H(\mu_{1})H^{*}(\mu_{1})]^{-1},\cdots,$
$\hskip 113.81102pt2\text{Real}(\mu_{w})[I_{p}+H(\mu_{w})H^{*}(\mu_{w})]^{-1}\big).$ (80)
For $D$-$\mathcal{H}_{\infty}$-BT:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\sigma_{1})[I_{m}+(1-\gamma^{-2})H^{*}(\sigma_{1})H(\sigma_{1})]^{-1},\cdots,$
$\hskip 85.35826pt2\text{Real}(\sigma_{v})[I_{m}+(1-\gamma^{-2})H^{*}(\sigma_{v})H(\sigma_{v})]^{-1}\big),$ (81)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\mu_{1})[I_{p}+(1-\gamma^{-2})H(\mu_{1})H^{*}(\mu_{1})]^{-1},\cdots,$
$\hskip 85.35826pt2\text{Real}(\mu_{w})[I_{p}+(1-\gamma^{-2})H(\mu_{w})H^{*}(\mu_{w})]^{-1}\big).$ (82)
For DD-PR-BT:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\sigma_{1})[I_{m}-t_{v,1}^{*}H^{*}(\sigma_{1})R_{\text{PR}}^{-1}H(\sigma_{1})t_{v,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\text{Real}(\sigma_{v})[I_{m}-t_{v,v}^{*}H^{*}(\sigma_{v})R_{\text{PR}}^{-1}H(\sigma_{v})t_{v,v}]^{-1}\big),$ (83)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\mu_{1})[I_{p}-t_{w,1}^{*}H(\mu_{1})R_{\text{PR}}^{-1}H^{*}(\mu_{1})t_{w,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\text{Real}(\mu_{w})[I_{p}-t_{w,w}^{*}H(\mu_{w})R_{\text{PR}}^{-1}H^{*}(\mu_{w})t_{w,w}]^{-1}\big),$ (84)
$T_{v}$ $\approx\text{blkdiag}\big(t_{v,1},\cdots,t_{v,v}\big),$ (85)
$T_{w}$ $\approx\text{blkdiag}\big(t_{w,1},\cdots,t_{v,w}\big),$ (86)

### v3

@@ 3.6 Numerical Issues and Their Solutions
4 Data-driven Projection-based Approximations of Various Generalizations of BT
ADI-based approximations of various generalizations of BT are non-intrusive but not data-driven, as samples of $G(s)$ in the right-half $s$-plane cannot be measured directly in an experimental setting. In ADI methods, the shifts must have a non-zero real part; this effectively makes it theoretically impossible to implement ADI-based approximations of various generalizations of BT using measurable data (i.e., samples of $G(s)$ on the imaginary axis). This theoretical limitation is reflected in the computation of the free parameter $\zeta$ within the interpolants $\hat{G}(s)=C\hat{V}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ and $\hat{G}(s)=\zeta(sI-S_{w}+L_{w}\zeta)^{-1}\hat{W}^{*}B+D$, because computing $\zeta$ requires the solution of Lyapunov equations, which only have unique solutions when the interpolation points are in the right-half of the $s$-plane.
Let $\mathcal{P}$ and $\mathcal{Q}$ denote the generalized controllability and observability Gramians that arise in various generalizations of BT. The key observation in [16] is that if
$\mathcal{P}\approx\hat{V}\hat{L}_{p}\hat{L}_{p}^{*}\hat{V}^{*},\qquad\mathcal{Q}\approx\hat{W}\hat{L}_{q}\hat{L}_{q}^{*}\hat{W}^{*},$
where $\hat{V}$ and $\hat{W}$ are defined in (1) and (2) with all interpolation points $\sigma_{i}$ and $\mu_{i}$ on the imaginary axis, and the factors $\hat{L}_{p}$ and $\hat{L}_{q}$ can be computed non-intrusively, then a data-driven implementation of any such generalization of BT follows directly. We note that if $\mathcal{P}$ and $\mathcal{Q}$ are approximated via Krylov subspace–based Petrov–Galerkin projection, approximations of the form $\mathcal{P}\approx\hat{V}\hat{L}_{p}\hat{L}_{p}^{*}\hat{V}^{*}$ and $\mathcal{Q}\approx\hat{W}\hat{L}_{q}\hat{L}_{q}^{*}\hat{W}^{*}$ arise naturally. The spectral factorizations appearing in [47] result from rewriting the Riccati equations in these BT variants as integral expressions so that they can be approximated by numerical integration. This manipulation leads to the need for samples of certain spectral factorizations, which cannot be measured in an experimental setting. Since ADI methods are special cases of Krylov subspace–based methods, as noted in the previous section, the corresponding non-intrusive implementations require only samples of $G(s)$. The requirement to sample in the right half of the $s$-plane stems from the theoretical conditions on the ADI shifts; it is not inherent to the Krylov subspace–based projection framework itself. We now present a proof-of-concept example to illustrate that the Gramians $\mathcal{P}$ and $\mathcal{Q}$ in the BT generalizations considered here can be approximated using interpolation points $\sigma_{i}$ and $\mu_{i}$ on the imaginary axis, while the factors $\hat{L}_{p}$ and $\hat{L}_{q}$ are still computed non-intrusively.
Illustrative Example
Consider an $8^{\text{th}}$-order LTI model with the following state-space realization:
$A$ $=\begin{bmatrix}-22.1414&-14.1915&-35.8543&-7.8301&54.2479&-1.6149&4.5713&-47.9895\\ -1.4098&6.0485&-2.0663&36.2832&88.6974&15.7929&74.5229&-30.3651\\ -20.9974&-3.8320&-7.5951&-40.6679&-71.4159&-45.2401&-39.4774&-4.9186\\ -107.6001&-67.2020&-108.3961&-33.5432&43.9644&-77.3467&21.1433&-109.2904\\ 143.7150&66.7452&146.1617&81.1110&-20.7467&135.0017&12.6135&158.4851\\ -101.2836&-48.3008&-93.7796&-47.9424&-20.5617&-89.5448&-28.7377&-78.5117\\ 129.4943&22.2846&56.8624&112.4968&85.9715&148.5772&61.3542&106.4938\\ 65.1568&63.9156&113.2290&-42.8674&-170.4269&-12.2243&-97.6582&98.1684\end{bmatrix}$,
$B$ $=\begin{bmatrix}0.6007&1.6263&-0.4206&2.5576&-2.1955&1.2682&-0.4758&-3.1936\end{bmatrix}^{T},$
$C$ $=\begin{bmatrix}1.9237&1.2498&1.3247&1.5407&1.1059&1.3546&0.9754&1.2928\end{bmatrix}$,
$D$ $=0.2378.$
This model is stable, minimum phase, square, and passive, with $\|G(s)\|_{\mathcal{H}_{\infty}}=0.4999$. Hence all generalizations of BT considered in this paper are applicable to this example.
For $\mathcal{H}_{\infty}$BT, set $\gamma=2$. Let the right interpolation points be $\sigma_{i}=(j9.99,-j9.99,j19.99,-j19.99,j29.99,-j29.99)$. Next, construct the interpolant $\hat{G}(s)=C\hat{V}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ from the samples $\big(G(j9.99),G(j19.99),G(j29.99),G(\infty)\big)$ by choosing the free parameter $\zeta$ as
$\zeta=\begin{bmatrix}1.0075+j0.0417\\ 1.0075-j0.0417\\ 1.0080-j0.0792\\ 1.0080+j0.0792\\ 0.9845-j0.2113\\ 0.9845+j0.2113\end{bmatrix}$.
@@ 4 Data-driven Projection-based Approximations of Various Generalizations of BT
Dually, construct the interpolant $\hat{G}(s)=\zeta(sI-S_{w}+L_{w}\zeta)^{-1}\hat{W}^{*}B+D$ from the samples $\big(G(j10),G(j20),G(j30),G(\infty)\big)$ by choosing
$\zeta=\begin{bmatrix}1.0075+j0.0417\\ 1.0075-j0.0417\\ 1.0080-j0.0792\\ 1.0080+j0.0792\\ 0.9845-j0.2113\\ 0.9845+j0.2113\end{bmatrix}^{T}.$
Next, solve the projected matrix equations by replacing $(A,B,C)$ with $(S_{w}-L_{w}\zeta,\hat{W}^{*}B,\zeta)$ in (15), (17), (19), (21), (23), and (24) to compute $\hat{Q}$, $\hat{Q}_{\mathrm{LQG}}$, $\hat{Q}_{\mathcal{H}_{\infty}}$, $\hat{Q}_{\mathrm{PR}}$, $\hat{Q}_{\mathrm{BR}}$, and $\hat{Q}_{\mathrm{SW}}$, respectively. Furthermore, compute $\hat{Q}_{\mathrm{S}}$ by replacing $(\tilde{A}_{\mathrm{S}},\tilde{B}_{\mathrm{S}},\tilde{C}_{\mathrm{S}})$ with $\big(S_{w}-L_{w}\zeta-\hat{W}^{*}BR_{\mathrm{S}}^{-1}\zeta,(\hat{W}^{*}\hat{V}\hat{P}\hat{V}^{*}C^{T}+\hat{W}^{*}BD^{T})R_{\mathrm{S}}^{-\frac{1}{2}},R_{\mathrm{S}}^{-\frac{1}{2}}\zeta\big)$ in (78).
These non-intrusively computed Gramians of the projected state-space realizations are factorized as $\hat{L}_{p}\hat{L}_{p}^{*}$ and $\hat{L}_{q}\hat{L}_{q}^{*}$. Thereafter, the respective Gramians $\mathcal{P}$ and $\mathcal{Q}$ are replaced with their Petrov–Galerkin projection–based approximations $\mathcal{P}\approx\hat{V}\hat{L}_{p}\hat{L}_{p}^{*}\hat{V}^{*}$ and $\mathcal{Q}\approx\hat{W}\hat{L}_{q}\hat{L}_{q}^{*}\hat{W}^{*}$, respectively, in BSA, leading to a non-intrusive, data-driven implementations of BT, LQGBT, $\mathcal{H}_{\infty}$BT, PRBT, BRBT, SWBT, and BST constructing the ROMs from the samples $\big(G(j9.99),G(j10),G(j19.99),G(j20),G(j29.99),G(j30),G(\infty)\big)$. The relative error $\frac{\|G(s)-\hat{G}(s)\|_{\mathcal{H}_{\infty}}}{\|G(s)\|_{\mathcal{H}_{\infty}}}$ for order-3 ROMs produced by the intrusive and non-intrusive versions of these BT variants is given in Table 1. The results show that the intrusive and non-intrusive implementations perform indistinguishably.
Table 1: Relative Error Comparison for the ROMs
Method $\frac{\|G(s)-\hat{G}(s)\|_{\mathcal{H}_{\infty}}}{\|G(s)\|_{\mathcal{H}_{\infty}}}$
Intrusive BT $0.4039$
Non-intrusive BT $0.4039$
Intrusive LQGBT $0.4037$
Non-intrusive LQGBT $0.4037$
Intrusive $\mathcal{H}_{\infty}$BT $0.4038$
Non-intrusive $\mathcal{H}_{\infty}$BT $0.4037$
Intrusive PRBT $0.4014$
Non-intrusive PRBT $0.4013$
Intrusive BRBT $0.4045$
Non-intrusive BRBT $0.4045$
Intrusive SWBT $0.4014$
Non-intrusive SWBT $0.4014$
Intrusive BST 0.4014
Non-intrusive BST 0.4014
The elegance of ADI methods lies in the fact that they provide a free parameter $\zeta$ that guarantees the projected Lyapunov equation admits a unique positive-definite solution and the projected Riccati equation admits a unique stabilizing positive-definite solution. This is achieved through their pole placement properties. In particular, LRCF-ADI places the poles of $\hat{A}$ at the conjugates of ADI shifts $\sigma_{i}$, while RADI places the poles of $\hat{A}-\hat{P}_{\mathrm{LQG}}\hat{C}^{*}\hat{C}$ at the conjugates of ADI shifts $\sigma_{i}$. Since the pole locations and the ADI shifts are tied together, the shifts $\sigma_{i}$ must be Hurwitz in LRCF-ADI and RADI. In a general setting, however, the interpolation points $\sigma_{i}$ and the poles of $\hat{A}$ and $\hat{A}-\hat{P}_{\mathrm{LQG}}\hat{C}^{*}\hat{C}$ need not be linked in Krylov subspace–based projection methods. Therefore, it is possible to choose the free parameter $\zeta$ so that the projected Lyapunov equation admits a unique positive-definite solution and the projected Riccati equation admits a unique stabilizing positive-definite solution even when the interpolation points $\sigma_{i}$ lie anywhere in the $s$-plane, including the $j\omega$-axis. In this section, we give specific choices of $\zeta$ that ensure the projected Lyapunov equation admits a unique positive-definite solution and the projected Riccati equation admits a unique stabilizing positive-definite solution when the interpolation points lie on the imaginary axis instead of in the right half of the $s$-plane.
4.1 Data-driven Projection-based Approximation of BT
Assume that all interpolation points $\sigma_{i}$ are located on the imaginary axis, i.e., $\sigma_{i}=j\omega_{i}$, so that the interpolant $\hat{G}(s)=\hat{C}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ can be computed from measurable data $G(j\omega_{i})$. Recall that the projected matrices satisfy the following:
$\hat{A}=S_{v}-\hat{B}L_{v},\quad\hat{B}=\zeta,\quad\hat{C}=\begin{bmatrix}H(j\omega_{1})&\cdots&H(j\omega_{v})\end{bmatrix},\quad D=G(\infty),$
where $S_{v}=\mathrm{diag}(j\omega_{1},\dots,j\omega_{v})\otimes I_{m}$ and $L_{v}=\begin{bmatrix}1&\cdots&1\end{bmatrix}\otimes I_{m}$.
@@ 4.1 Data-driven Projection-based Approximation of BT
Proposition 11.
Let $(\sigma_{1},\dots,\sigma_{v})$ be $v$ distinct interpolation points, and let $(\lambda_{1},\dots,\lambda_{v})$ be $v$ distinct desired poles such that the two sets have no elements in common. Define
$S_{p}=\mathrm{diag}(-\overline{\lambda_{1}},\dots,-\overline{\lambda_{v}})\otimes I_{m}.$
Further, assume that the pairs $(-S_{v},L_{v})$ and $(-S_{p},L_{v})$ are observable. Let $X_{p}$ be the unique solution of the Sylvester equation
$-S_{p}^{*}X_{p}-X_{p}S_{v}+L_{v}^{T}L_{v}=0.$ (116)
Then the interpolant
$\hat{A}=S_{v}-\hat{B}L_{v},\quad\hat{B}=X_{p}^{-1}L_{v}^{T},\quad\hat{C}=\begin{bmatrix}H(\sigma_{1})&\cdots&H(\sigma_{v})\end{bmatrix},\quad D=G(\infty)$ (117)
satisfies the interpolation condition $G(\sigma_{i})=\hat{G}(\sigma_{i})$ and has poles at $(\lambda_{1},\dots,\lambda_{v})$ with multiplicity $m$.
Proof.
The assumption that the sets $(\sigma_{1},\dots,\sigma_{v})$ and $(\lambda_{1},\dots,\lambda_{v})$ are disjoint ensures the uniqueness of the Sylvester equation (116). Moreover, the observability of the pairs $(-S_{v},L_{v})$ and $(-S_{p},L_{v})$ guarantees that $X_{p}$ is invertible. Premultiplying (116) by $X_{p}^{-1}$ gives
$-X_{p}^{-1}S_{p}^{*}X_{p}-S_{v}+X_{p}^{-1}L_{v}^{T}L_{v}=0,$
$S_{v}-\hat{B}L_{v}=-X_{p}^{-1}S_{p}^{*}X_{p},$
$\hat{A}=-X_{p}^{-1}S_{p}^{*}X_{p}.$
Hence, the eigenvalues of $\hat{A}$ coincide with those of $S_{p}^{*}$, and the ROM (117) has the desired poles $(\lambda_{1},\dots,\lambda_{v})$ with multiplicity $m$. ∎
The next proposition focuses on ensuring the invertibility of the matrix $X_{p}$ when the interpolation points $\sigma_{i}$ are located on the imaginary axis.
Proposition 12.
Assume that Proposition 11 holds. Further, for $\epsilon>0$, let $\sigma_{i}=j\omega_{i}$ and $\lambda_{i}=-\epsilon+j\omega_{i}$, satisfying
$\Delta_{\min}:=\min_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{v}|\omega_{i}-\omega_{j}|>0.$
Then $X_{p}$ admits the explicit Kronecker form $X_{p}=Y\otimes I_{m}$, where $Y\in\mathbb{C}^{v\times v}$ has entries
$Y_{ij}=\frac{1}{\epsilon+j(\omega_{j}-\omega_{i})},\qquad i,j=1,\dots,v.$ (118)
The following properties hold:
  1. 1.
    $X_{p}$ is asymptotically diagonal as $\epsilon\to 0$ with $X_{p}=\frac{1}{\epsilon}I_{vm}+\mathcal{O}(1)$.
  2. 2.
    By defining the diagonal approximation $\widetilde{X}_{p}:=\frac{1}{\epsilon}I_{vm}$, the relative error in Frobenius norm satisfies
    $\frac{\|X_{p}-\widetilde{X}_{p}\|_{F}}{\|\widetilde{X}_{p}\|_{F}}\leq\frac{\epsilon\sqrt{v-1}}{\Delta_{\min}}.$ (119)
    In particular, for any tolerance $\tau\in(0,1)$, if
    $\epsilon<\tau\,\frac{\Delta_{\min}}{\sqrt{v-1}},$ (120)
    then $\|X_{p}-\widetilde{X}_{p}\|_{F}/\|\widetilde{X}_{p}\|_{F}<\tau$.
  3. 3.
    $X_{p}$ is strictly diagonally dominant whenever
    $\epsilon<\frac{\Delta_{\min}}{v-1}.$ (121)
Proof.
The proof is given in Appendix D. ∎
By selecting a small positive scalar $\epsilon$ as in Proposition 12, we can effectively impose a modal structure with lightly damped stable poles on the ROM (117); that is, as $\epsilon\to 0$,
$\hat{A}\approx\mathrm{diag}(-\epsilon+j\omega_{1},\cdots,-\epsilon+j\omega_{v})\otimes I_{m},\quad\hat{B}\approx\begin{bmatrix}\epsilon&\cdots&\epsilon\end{bmatrix}^{T}\otimes I_{m}.$
[…]
Proposition 13.
Assume that Proposition 12 holds. Let $\hat{A}\in\mathbb{C}^{vm\times vm}$ and $\hat{B}\in\mathbb{C}^{vm\times m}$ be defined as:
$\hat{A}=S_{v}-\hat{B}L_{v},\quad\hat{B}=X_{p}^{-1}L_{v}^{T},$
[…]
$\hat{A}\hat{P}+\hat{P}\hat{A}^{*}+\hat{B}\hat{B}^{*}=0.$ (122)
For any tolerance $\delta\in(0,1)$, if $\epsilon$ satisfies
$0<\epsilon\leq\min\left(\frac{\Delta_{\min}}{4v},\quad\frac{\delta\cdot\Delta_{\min}}{8v^{2}}\right),$ (123)
then the following properties hold:
  1. 1.
    $\hat{A}$ is Hurwitz, and $\hat{P}$ is Hermitian positive definite.
  2. 2.
    $\hat{P}$ is strictly block diagonally dominant. Specifically, let $\hat{P}_{d}=\mathrm{blkdiag}(\hat{P}_{11},\dots,\hat{P}_{vv})$ denote the block diagonal part of $\hat{P}$. Then for every block row $i$:
    $\sum_{j\neq i}\|\hat{P}_{ij}\|\leq\delta\cdot\sigma_{\min}(\hat{P}_{ii}).$
  3. 3.
    The diagonal blocks satisfy the asymptotic limit:
    $\lim_{\epsilon\to 0}\frac{1}{\epsilon}\hat{P}_{ii}=\frac{1}{2}I_{m},\qquad i=1,\dots,v.$
    Consequently, $\hat{P}=\frac{\epsilon}{2}I_{vm}+\mathcal{O}(\epsilon^{3})$.
  4. 4.
    The block diagonal approximation error in the induced infinity norm satisfies:
    $\|\hat{P}-\hat{P}_{d}\|_{\infty}\leq\delta\cdot\min_{i}\sigma_{\min}(\hat{P}_{ii}).$
Proof.
The proof is given in Appendix E. ∎
[…]
4.2 Data-driven Projection-based Approximation of LQGBT
In this subsection, similar to NI-ADI-LQG, $P_{\mathrm{LQG}}$ and $Q_{\mathrm{LQG}}$ are approximated via Petrov–Galerkin projection, i.e., $P_{\mathrm{LQG}}\approx\hat{V}\hat{P}_{\mathrm{LQG}}\hat{V}^{*}$ and $Q_{\mathrm{LQG}}\approx\hat{W}\hat{Q}_{\mathrm{LQG}}\hat{W}^{*}$. To this end, we need to choose the free parameter $\zeta$ such that the projected Riccati equations (29) and (32) admit positive-definite stabilizing solutions. Theorems 1 and 2 provide specific choices of $\zeta$ that achieve this when the interpolation points lie in the right half of the $s$-plane. The focus here is to develop a method to compute $\zeta$ that ensures this property when the interpolation points lie on the imaginary axis.
Theorem 14.
Consider the algebraic Riccati equation (ARE)
$\hat{A}(\epsilon)\hat{P}_{\mathrm{LQG}}+\hat{P}_{\mathrm{LQG}}\hat{A}(\epsilon)^{*}+\hat{B}(\epsilon)\hat{B}(\epsilon)^{*}-\hat{P}_{\mathrm{LQG}}\hat{C}^{*}\hat{C}\hat{P}_{\mathrm{LQG}}=0,$ (124)
where $\epsilon>0$, and the system matrices are defined as:
$\hat{A}(\epsilon)$ $=\mathrm{diag}(-\epsilon+j\omega_{1},\cdots,-\epsilon+j\omega_{v})\otimes I_{m},$
$\hat{B}(\epsilon)$ $=\begin{bmatrix}\epsilon I_{m}\\ \vdots\\ \epsilon I_{m}\end{bmatrix}$,
$\hat{C}$ $=\begin{bmatrix}H(j\omega_{1})&\cdots&H(j\omega_{v})\end{bmatrix}$.
Assume the non-zero frequencies $\omega_{1},\dots,\omega_{v}\in\mathbb{R}$ are distinct, and let $\Delta_{\mathrm{min}}:=\min_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{v}|\omega_{i}-\omega_{j}|>0$. Assume further that each $H(j\omega_{i})$ has full column rank. Then, there exist constants $\epsilon_{0}>0$ and $K>0$ such that for all $0<\epsilon\leq\epsilon_{0}$, the unique stabilizing solution $\hat{P}_{\mathrm{LQG}}(\epsilon)$ exists and admits the decomposition:
$\hat{P}_{\mathrm{LQG}}(\epsilon)=\tilde{P}_{\mathrm{LQG}}(\epsilon)+E(\epsilon),$ (125)
where $\tilde{P}_{\mathrm{LQG}}(\epsilon)=\mathrm{blkdiag}(p_{1}(\epsilon),\dots,p_{v}(\epsilon))$ with
$p_{i}(\epsilon)=\epsilon\big(H^{*}(j\omega_{i})H(j\omega_{i})\big)^{-1}\Big(\big(I_{m}+H^{*}(j\omega_{i})H(j\omega_{i})\big)^{\frac{1}{2}}-I_{m}\Big),$ (126)
and the error term satisfies $\|E(\epsilon)\|\leq K\epsilon^{2}$.
Proof.
The proof is given in Appendix F. ∎
Theorem 14 indicates that the projected Riccati equation (29) has a stabilizing solution when the projected matrix $\hat{A}$ is Hurwitz and has a modal form with lightly damped poles, which is consistent with the observation made in [13, 14]. Next, it is shown how to enforce this structure on $\hat{A}$ using Proposition 11.
Proposition 15.
Let $\epsilon>0$ and let $\omega_{1},\dots,\omega_{v}\in\mathbb{R}$ be distinct non-zero frequencies. Define the minimum frequency magnitude $\omega_{\min}:=\min_{i}|\omega_{i}|$ and the minimum frequency separation $\Delta_{\mathrm{min}}:=\min_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{v}|\omega_{i}-\omega_{j}|>0$. Consider the matrix $\hat{A}\in\mathbb{C}^{vm\times vm}$ defined by
$\hat{A}=S_{v}-\hat{B}L_{v},$
where $S_{v}=\mathrm{diag}(j\omega_{1},\dots,j\omega_{v})\otimes I_{m}$, $L_{v}=(\mathbf{1}_{v}^{T}\otimes I_{m})$, and $\hat{B}=(\epsilon\mathbf{1}_{v}\otimes I_{m})$, with $\mathbf{1}_{v}\in\mathbb{R}^{v}$ denoting the vector of all ones. Assume $v\geq 2$. For any admissible relative error tolerance $\delta\in(0,1)$, if $\epsilon$ satisfies
$0<\epsilon\leq\min\left(\frac{\delta\cdot\omega_{\min}}{v-1},\quad\frac{\Delta_{\mathrm{min}}}{2(v-1)}\right),$ (127)
then the following properties hold:
  1. 1.
    $\hat{A}$ is strictly diagonally dominant by rows.
  2. 2.
[…]
    $\|\hat{A}-\hat{A}_{d}\|_{\infty}\leq\delta\cdot\omega_{\min}.$
  3. 3.
    The eigenvalues of $\hat{A}$ are distinct and lie within $v$ disjoint Gershgorin disks in the complex plane, each centered at $\epsilon+j\omega_{i}$ with radius $(v-1)\epsilon$.
Proof.
The proof is given in Appendix G. ∎
Thus, by selecting $\epsilon$ according to Theorem 14 and Proposition 15, one can obtain the free parameter $\hat{B}$ that ensures the projected Riccati equation (29) admits a positive-definite stabilizing solution. Dually, the free parameter $\hat{C}$ can be obtained to ensure that the projected Riccati equation (32) admits a positive-definite stabilizing solution. Consequently, $P_{\mathrm{LQG}}$ and $Q_{\mathrm{LQG}}$ admit projection-based approximations
$P_{\mathrm{LQG}}\approx\hat{V}\hat{P}_{\mathrm{LQG}}\hat{V}^{*}=\hat{V}\hat{L}_{p}\hat{L}_{p}^{*}\hat{V}^{*},\quad Q_{\mathrm{LQG}}\approx\hat{W}\hat{Q}_{\mathrm{LQG}}\hat{W}^{*}=\hat{W}\hat{L}_{q}\hat{L}_{q}^{*}\hat{W}^{*}.$
Replacing the exact Gramian factors $L_{p}$ and $L_{q}$ in BSA with the approximations $\tilde{L}_{p}=\hat{V}\hat{L}_{p}$ and $\tilde{L}_{q}=\hat{W}\hat{L}_{q}$ yields a data-driven projection-based approximation of LQGBT (DD-P-LQGBT). This approach enables the implementation of DD-P-LQGBT using only measurable data $G(j\omega_{i})$, $G(j\nu_{i})$, and $G(\infty)$.
Furthermore, if the amount of data makes it computationally infeasible to solve the projected Riccati equations (29) and (32), one can use diagonally dominant approximations of $\hat{P}_{\mathrm{LQG}}$ and $\hat{Q}_{\mathrm{LQG}}$ from Theorem 14 as follows:
$\hat{P}_{\mathrm{LQG}}$ $\approx\tilde{P}_{\mathrm{LQG}}=\mathrm{blkdiag}(p_{1},\dots,p_{v}),$ (128)
$\hat{Q}_{\mathrm{LQG}}$ $\approx\tilde{Q}_{\mathrm{LQG}}=\mathrm{blkdiag}\big(q_{1},\cdots,q_{w}\big),$ (129)